OpenStudy (anonymous):

Write the slope-intercept form of the equation for the line passing through the pint (-6, -5) and parallel to the line y=4x+6

OpenStudy (jdoe0001):

hmm what's the slope of " y=4x+6 " ?

OpenStudy (anonymous):

4/1

OpenStudy (jdoe0001):

right.. 4/1 or just 4, and any parallel line to that one, will also have the same slope as thata one so you'd want the equation of a line whose slope is 4 and passes through (-6, -5), or \(\bf \begin{array}{lllll} &x_1&y_1\\ &({\color{red}{ -6}}\quad ,&{\color{blue}{ -5}})\quad \end{array} \\\quad \\ slope = {\color{green}{ m}}= 4 \\ \quad \\ y-{\color{blue}{ y_1}}={\color{green}{ m}}(x-{\color{red}{ x_1}})\qquad \textit{plug in the values and solve for "y"}\\ \qquad \uparrow\\ \textit{point-slope form} \)

OpenStudy (anonymous):

So y-(-5)=4(x-(-6))?

OpenStudy (jdoe0001):

yes, that'd give you the equation in point-slope form but you're asked to give the slope-intercept form, so solving that for "y" will give you the slope-intercept form